g(x) = 2x -1, 8)) Given f(x) = x?, a) f(g(x)) h(x) = Vx+2; find the following: b) g(h(x))
Answer #1

-F ()=22 9(W) = 2x-1 h(@DEVX+2 - a) Flaca)) = 182-12 = 22213+ (0)2 -2X22 = 422 +1-48 = 140²-48+ (b) g (h(W) V 22-1+2 /22+1 1/